AlexisR:

how do i do this correctly my profesor sed im missing sum steps and why i changed the 2 to a 1 i had help from a previus person but i dont know what i did wron 2*│x-4│=2 2*│x-4│=1 2*│x-4│=-1 2x=1+4 2x=-1+4 x=5/2 x=3/2

mhanifa:

Is the original equation 2|x - 4| = 2?

mhanifa:

If so the first step is to divide both sides by 2.

AlexisR:

yes 2|x - 4| = 2 is the first one

mhanifa:

Can you do the first step?

AlexisR:

do i devide |x - 4| by 2?

mhanifa:

divide 2|x - 4| by 2, left side and divide 2 by 2, right side

mhanifa:

What you get?

AlexisR:

2/x-1/2=1

mhanifa:

2|x - 4| / 2 = |x - 4| 2/2 = 1

AlexisR:

i think i keep the | right?

mhanifa:

You now have: |x - 4| = 1

AlexisR:

i see

mhanifa:

Can you continue from here?

mhanifa:

Anything unclear at this step?

AlexisR:

this is the next step right? │x-4│=1 │x-4│=-1

mhanifa:

Right, its correct

mhanifa:

But without absolute value symbols

mhanifa:

x - 4 = 1 x - 4 = -1

mhanifa:

Can you solve it from here?

AlexisR:

i see ok ok

mhanifa:

What is your solutions then?

AlexisR:

x-4=1 x-4=-1 x=1+4 x=1+4 X=5/1 X=3/1

AlexisR:

right?

mhanifa:

Why you put x = 5/1, it should be x = 5, same with the other one. Otherwise all good, well done!

AlexisR:

oh yah i grabd the 1 from the x basecly 1x but now that you day that yah it is done!

AlexisR:

waw thx its correct

AlexisR:

i can finally get my highschool degree after completing this last thing

mhanifa:

Just to recap I'll add full solution here: 2|x - 4| = 2 |x - 4| = 1 x - 4 = 1 x - 4 = - 1 x = 1 + 4 x = -1 + 4 x = 5 x = 3 Thanks and good luck with your assignment!

AlexisR:

u too have a nice day